OpenStudy (anonymous):

Solve the equation: (3x - 2)(7x + 1) = 0. A. {2/3, -1/7} B. {-2/3, 1/7} C. 2/3 only D. -1/7 only E. no solution

OpenStudy (anonymous):

C. 2/3 only

OpenStudy (apoorvk):

thsi is a pair of equations  bla- bla part and clop clop parts will always equal zero. therefore 3x-2=0 ; so x= 2/3. and also, 7x+1=0 ; so x= -1/7 hence answer is [A.]